PHOTOELECTRON AND INFRARED SPECTRA OF (H_{2}O)_{n}^{−} CLUSTERS: NEW ISOMERS AND THE SPECTRAL IDENTIFICATION OF ELECTRON BINDING MOTIFS

\underline{JOSEPH R. ROSCIOLI}, JOSEPH C. BOPP, NATHAN I. HAMMER, MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, PO Box 208107, New Haven, CT 06520.

We present infrared and photoelectron spectra of the argon-solvated water hexamer and heptamer anions. The photoelectron spectra of both species are dominated by two prevalent isomer classes, I and II. However, the heptamer develops a new, high-binding isomer class when more than seven argon atoms are attached. We also present recently obtained HOH bending spectra of large, warm water cluster anions, and find a prominent transition corresponding to a double-acceptor (AA) water molecule. This suggests the presence of the AA binding motif in isomer I into the large size regime.